2. For each of the following, find the market equilibrium price and quantity (a) Let Q,-3P-20 and Qa 40-2P. (b) Let Q, -6P and Qa 40 2P. 3. For each of the following, draw the effects of the shocks on a supply and demand graph. What happens to the equilibrium price and quantity in each of the scenarios? (a) Suppose that, due to the rising popularity of Moscow Mules, demand for copper mugs his risen. At the same time, due to global trade restrictions, the price of raw copper has increased. (b) Suppose that the country has fallen into a recession. As a result, household incomes have fallen. This has caused demand for TVs, a normal good, to decrease. At the same time, several TV manufacturers have gone out of business.
Answer #1

Q2. a) Qs = 3P - 20, Qd = 40 - 2P

Market equillibrium price and quantity

Market equillibrium is where the demand and supply match each other.

So, 3P - 20 = 40 - 2P

3P + 2P = 40 + 20

5P = 60

P = 12

Q = 3*12 - 20

= 36 - 20

= 16

b) Qs = 6P and Qd = 40 - 2P

6P = 40 - 2P

6P + 2P = 40

8P = 40

P = 5

Q = 6*5

= 30

Q3. a) As the demand for copper mugs has risen, the demand curve shifts rightwards. At the same time, as the price of raw copper increases, so the supply of copper mugs would come down and the supply curve shifts leftwards. Thus the new equillibrium price rises and the equillibrium quantity would remain more or less the same.

b) As the demand for TV's decreased, the demand curve shifts leftwards. Moreover, as several TV manufacturers have gone out of business, the supply would come down and the supply curve shifts leftwards. Thus the new equillibrium quantity reduces but the equillibrium price remains more or less the same.